# Cell-cell communication inference and analysis: biological mechanisms, computational approaches, and future opportunities

### v1

@@ 6.3 De novo construction of CCC
De novo construction of cellular signaling networks represents an important frontier to infer CCC directly from single-cell and spatial omics data without reliance on prior knowledge of ligand-receptor pairs. This approach offers potential to uncover novel signaling pathways and intercellular interactions mediated by complex mechanisms, thereby providing fresh biological insights. CytoTalk[RN318] pioneered this paradigm by reconstructing signaling pathways from scRNA-seq data without using known pathway annotations. Its algorithm formulates this as a prize-collecting Steiner forest (PCSF) problem, identifying context-specific, parsimonious subnetworks that connect salient genes within and between cell types. The method operates on the rationale that cell-type-specific signals are encoded in highly expressed genes that are topologically proximate to active ligand-receptor pairs within a background protein-protein interaction network. While this approach reveals context-specific signaling cascades absent from canonical pathways, its utility is constrained by the completeness of the background network and computational demands for large datasets. For spatial transcriptomics with limited gene panels and inadequate ligand-receptor coverage, de novo construction offers distinct advantages. Methods including DeepLinc[RN362DeepLinc], Spacia[RN432Spacia], GITIII[Xiao20240821608964GITIII], AMICI[RN456AMICI] and Steamboat[LiangSteamboat] infer CCC by integrating gene expression with spatial proximity without preselecting genes based on prior knowledge. However, omitting ligand-receptor information presents a fundamental challenge: establishing causal relationships for predicted interactions and distinguishing whether associated genes function as signal mediators or downstream targets.
@@ 6.4 Mapping CCC across space, time and conditions
Spatial context is crucial for accurate CCC inference (Figure 7c), yet current spatial transcriptomics technologies vary substantially in cellular resolution, gene coverage, and signal-to-noise ratio, creating significant challenges for single-cell resolution analysis. Low-resolution methods like Visium[staahl2016visualization] capture mixed cell populations within individual spots, complicating heterogeneity studies, while high-resolution in situ techniques such as Xenium and CosMx[he2022high] offer single-cell resolution but suffer from limited transcript detection. The large scale, technical noise, and sparsity of high-resolution spatial data impose stringent requirements on CCC method design.
Temporal dynamics represent another critical dimension, enabling elucidation of communication mechanism evolution. Common approaches integrate cell lineage trajectory analysis with CCC inference to identify stage-specific LRIs along developmental trajectories[RN282, RN292], while other models leverage multi-timepoint data to capture interaction dynamics[RN413]. However, transcriptomic data alone provides limited temporal resolution, as ligand and receptor expression may not coincide temporally with functional communication events, underscoring the need for multi-omics integration.
CCC events occur across diverse temporal scales, with different processes—ligand-receptor binding, signal transduction, transcription factor activation, and downstream responses—operating at distinct rates that reflect biological function and pathway specificity. Paracrine signaling, relying on molecular diffusion, proceeds slowly: a typical protein requires hours to diffuse 1 millimeter in aqueous environments. In contrast, endocrine signaling molecules transported via fluid flow in the circulatory system can cover the same distance in approximately 3 seconds[muller2011extracellular]. Synaptic signaling operates even faster, with neuronal electrical impulses exceeding 100 m/s and neurotransmitter diffusion across synaptic clefts occurring within milliseconds[purves2017neuroscience]. Receptor activation is typically rapid—G protein-coupled receptors activate within milliseconds to microseconds[gruteser2022examination]. Cellular responses also vary considerably: protein modification-based responses (cell motility, secretion, metabolism) occur within seconds to minutes, while gene expression changes requiring new protein synthesis generally demand minutes to hours[alberts2022molecular]. Integrating dynamic equations that capture these temporal behaviors would enable more precise modeling of communication dynamics and enhance biological interpretation.
@@ 6.5 Linking CCC to phenotypes and clinical information
Disease progression involves dynamic tissue alterations and complex clinical manifestations. Integrating phenotypic and clinical data with CCC analysis can reveal interactions that play crucial roles in disease initiation and progression, offering new opportunities for phenotype manipulation, prognostic biomarker discovery, and therapeutic targeting[CCCreviewSu2024]. Such integration is essential for elucidating disease mechanisms and advancing precision medicine (Figure 7d). While some studies have identified phenotype-associated signaling pathways[RN425PathFinder] or used ligand–receptor interactions for prognosis prediction[yuan2019systematic], a key challenge lies in the heterogeneity of phenotypic and clinical data, and how to effectively integrate such data with molecular-level communication information within computational models. Furthermore, since phenotypic changes often arise from coordinated activity across multiple pathways—and vice versa—methods capable of disentangling these complex regulatory relationships are needed.
@@ 6.6 Combining CCC analysis with mechanism-driven modeling
Mathematical modeling enables exploration of biological processes across parameter ranges often inaccessible experimentally, facilitating evaluation of system dynamics under diverse conditions[daneshpour2019modeling] (Figure 7e). For instance, modeling revealed how IL-2’s dual role in T-cell proliferation and apoptosis contributes to immune homeostasis[hart2014paradoxical]. Combining mechanistic models with CCC analysis can help interpret inferred signaling patterns, validate pathway functionality, and uncover how signaling dynamics influence system homeostasis or disease progression[myers2021mechanistic]. This approach requires accurate representation of biological mechanisms and dynamic relationships, along with appropriate model assumptions—challenges that intensify in multi-scale systems involving diverse spatial and temporal dimensions.
Traditional mathematical modeling is typically limited to few cell types and pathways. Integrating genome-wide CCC analysis with mechanistic modeling remains challenging, though recent advances in learning cellular dynamics from single-cell or spatial transcriptomics show promise[zhang2025integrating]. Current methods for inferring RNA velocity or reconstructing vector fields from snapshot data, however, often overlook intercellular communication. Emerging approaches address this gap: one study incorporated a transformer module to capture CCC while learning cellular vector fields[jiang2025learning], and another developed CytoBridge, an unbalanced mean-field Schrödinger bridge method, to model cellular dynamics and interactions from time-series single-cell data[zhang2025modelingcelldynamicsinteractions, zhang2025deciphering]. These examples use neural networks to learn CCC without prior ligand–receptor information. Further refinement to incorporate communication more explicitly will not only improve the accuracy of the learned cellular dynamics but also offer new insights into temporal CCC dynamics and cellular responses to altered cellular communication.
@@ 6.7 Engineering CCC
Rapid advances in single-cell omics and artificial intelligence are fostering goal-directed, intervention-based approaches in CCC research, enabling deeper mechanistic insight and improved prediction. Key emerging directions include in silico perturbation, foundation model–based analysis, and virtual cell modeling (Figure 7f). In silico perturbation allows rapid assessment of genetic or chemical interventions on signaling and cell behavior, overcoming experimental limitations. For example, CellAgentChat uses an agent-based model to simulate perturbations via rule modifications, supporting novel intervention design[RN324CellAgentChat]. In addition, foundation models pretrained on large-scale single-cell and spatial datasets excel at tasks like cell annotation, multi-omic integration, and perturbation prediction[RN339LLM, hao2024large], yet adapting them for CCC analysis remains difficult due to ground-truth scarcity and pathway complexity. Virtual cells—computational models simulating cellular behavior and interactions—can leverage AI to replicate molecular, cellular, and tissue dynamics across conditions[bunne2024build]. However, robust evaluation frameworks and attention to interpretability are essential to ensure these methods reliably address biological questions.
@@ Acknowledgments
The work was supported by a Major Research Plan of the National Natural Science Foundation of China (no. 92374108). The authors thank Dr. Peijie Zhou in Peking University and Dr. Axel Almet in University of California, Irvine for reading the manuscript and helpful suggestions.

### v3

@@ 6.3 De novo construction of CCC
De novo construction of cellular signaling networks represents an important frontier to infer CCC directly from single-cell and spatial omics data without reliance on prior knowledge of ligand-receptor pairs. This approach offers potential to uncover novel signaling pathways and intercellular interactions mediated by complex mechanisms, thereby providing fresh biological insights. CytoTalk[79] pioneered this paradigm by reconstructing signaling pathways from scRNA-seq data without using known pathway annotations. Its algorithm formulates this as a prize-collecting Steiner forest (PCSF) problem, identifying context-specific, parsimonious subnetworks that connect salient genes within and between cell types. The method operates on the rationale that cell-type-specific signals are encoded in highly expressed genes that are topologically proximate to active ligand-receptor pairs within a background protein-protein interaction network. While this approach reveals context-specific signaling cascades absent from canonical pathways, its utility is constrained by the completeness of the background network and computational demands for large datasets. For spatial transcriptomics with limited gene panels and inadequate ligand-receptor coverage, de novo construction offers distinct advantages. Methods including DeepLinc[105], Spacia[210], GITIII[187], AMICI[75] and Steamboat[108] infer CCC by integrating gene expression with spatial proximity without preselecting genes based on prior knowledge. However, omitting ligand-receptor information presents a fundamental challenge: establishing causal relationships for predicted interactions and distinguishing whether associated genes function as signal mediators or downstream targets.
@@ 6.4 Mapping CCC across space, time and conditions
Spatial context is crucial for accurate CCC inference (Figure 7c), yet current spatial transcriptomics technologies vary substantially in cellular resolution, gene coverage, and signal-to-noise ratio, creating significant challenges for single-cell resolution analysis. Low-resolution methods like Visium[158] capture mixed cell populations within individual spots, complicating heterogeneity studies, while high-resolution in situ techniques such as Xenium and CosMx[72] offer single-cell resolution but suffer from limited transcript detection. The large scale, technical noise, and sparsity of high-resolution spatial data impose stringent requirements on CCC method design.
Temporal dynamics represent another critical dimension, enabling elucidation of communication mechanism evolution. Common approaches integrate cell lineage trajectory analysis with CCC inference to identify stage-specific LRIs along developmental trajectories[177, 101], while other models leverage multi-timepoint data to capture interaction dynamics[184]. However, transcriptomic data alone provides limited temporal resolution, as ligand and receptor expression may not coincide temporally with functional communication events, underscoring the need for multi-omics integration.
CCC events occur across diverse temporal scales, with different processes—ligand-receptor binding, signal transduction, transcription factor activation, and downstream responses—operating at distinct rates that reflect biological function and pathway specificity. Paracrine signaling, relying on molecular diffusion, proceeds slowly: a typical protein requires hours to diffuse 1 millimeter in aqueous environments. In contrast, endocrine signaling molecules transported via fluid flow in the circulatory system can cover the same distance in approximately 3 seconds[127]. Synaptic signaling operates even faster, with neuronal electrical impulses exceeding 100 m/s and neurotransmitter diffusion across synaptic clefts occurring within milliseconds[143]. Receptor activation is typically rapid—G protein-coupled receptors activate within milliseconds to microseconds[65]. Cellular responses also vary considerably: protein modification-based responses (cell motility, secretion, metabolism) occur within seconds to minutes, while gene expression changes requiring new protein synthesis generally demand minutes to hours[3]. Integrating dynamic equations that capture these temporal behaviors would enable more precise modeling of communication dynamics and enhance biological interpretation.
@@ 6.5 Linking CCC to phenotypes and clinical information
Disease progression involves dynamic tissue alterations and complex clinical manifestations. Integrating phenotypic and clinical data with CCC analysis can reveal interactions that play crucial roles in disease initiation and progression, offering new opportunities for phenotype manipulation, prognostic biomarker discovery, and therapeutic targeting[160]. Such integration is essential for elucidating disease mechanisms and advancing precision medicine (Figure 7d). While some studies have identified phenotype-associated signaling pathways[54] or used ligand–receptor interactions for prognosis prediction[195], a key challenge lies in the heterogeneity of phenotypic and clinical data, and how to effectively integrate such data with molecular-level communication information within computational models. Furthermore, since phenotypic changes often arise from coordinated activity across multiple pathways—and vice versa—methods capable of disentangling these complex regulatory relationships are needed.
@@ 6.6 Combining CCC analysis with mechanism-driven modeling
Mathematical modeling enables exploration of biological processes across parameter ranges often inaccessible experimentally, facilitating evaluation of system dynamics under diverse conditions[44] (Figure 7e). For instance, modeling revealed how IL-2’s dual role in T-cell proliferation and apoptosis contributes to immune homeostasis[69]. Combining mechanistic models with CCC analysis can help interpret inferred signaling patterns, validate pathway functionality, and uncover how signaling dynamics influence system homeostasis or disease progression[128]. This approach requires accurate representation of biological mechanisms and dynamic relationships, along with appropriate model assumptions—challenges that intensify in multi-scale systems involving diverse spatial and temporal dimensions.
Traditional mathematical modeling is typically limited to few cell types and pathways. Integrating genome-wide CCC analysis with mechanistic modeling remains challenging, though recent advances in learning cellular dynamics from single-cell or spatial transcriptomics show promise[203]. Current methods for inferring RNA velocity or reconstructing vector fields from snapshot data, however, often overlook intercellular communication. Emerging approaches address this gap: one study incorporated a transformer module to capture CCC while learning cellular vector fields[86], and another developed CytoBridge, an unbalanced mean-field Schrödinger bridge method, to model cellular dynamics and interactions from time-series single-cell data[204, 205]. These examples use neural networks to learn CCC without prior ligand–receptor information. Further refinement to incorporate communication more explicitly will not only improve the accuracy of the learned cellular dynamics but also offer new insights into temporal CCC dynamics and cellular responses to altered cellular communication.
@@ 6.7 Engineering CCC
Rapid advances in single-cell omics and artificial intelligence are fostering goal-directed, intervention-based approaches in CCC research, enabling deeper mechanistic insight and improved prediction. Key emerging directions include in silico perturbation, foundation model–based analysis, and virtual cell modeling (Figure 7f). In silico perturbation allows rapid assessment of genetic or chemical interventions on signaling and cell behavior, overcoming experimental limitations. For example, CellAgentChat uses an agent-based model to simulate perturbations via rule modifications, supporting novel intervention design[145]. In addition, foundation models pretrained on large-scale single-cell and spatial datasets excel at tasks like cell annotation, multi-omic integration, and perturbation prediction[41, 67], yet adapting them for CCC analysis remains difficult due to ground-truth scarcity and pathway complexity. Virtual cells—computational models simulating cellular behavior and interactions—can leverage AI to replicate molecular, cellular, and tissue dynamics across conditions[25]. However, robust evaluation frameworks and attention to interpretability are essential to ensure these methods reliably address biological questions.
@@ Acknowledgments
The authors thank Dr. Peijie Zhou in Peking University and Dr. Axel Almet in University of California, Irvine for reading the manuscript and helpful suggestions.
The work was supported by a Major Research Plan of the National Natural Science Foundation of China (Grant No. 92374108).
X. Cheng and H. Huang contributed equally to this work.